HOME ESR EXPERIMENTS LABS etc
NucAR lab laser_lab Fritz-Bosch-Lab X-Ray Lab target Beavertail
  Beavertail  Not logged in ELOG logo
Message ID: 43     Entry time: Sat Apr 5 14:11:55 2025
Author: CG, GL, ID 
Category: General 
Subject: Run stopped - detectors toast? 
run0034 stopped at 14:11

Leakage currents are increasing and once again we are not fully depleting the pad detectors at 150V.
Noticed that the strip vs energy histogram began with a well-defined band across all strips and the n-side for a given pad, but by the time we stopped event energies were spread up to a max value across all 
strips and n-side. Also saw *incredibly* long rise times and a rollover of the pre-amp pulse rather than a peak and decay.
  -> This implies loss of bias to the detector.

We increased the bias voltage to 300V on Bus 0, Mod 1, Ch 2 (6th SiPad on the inside pocket) and saw the leakage current rise in tandem with the voltage (up to 15uA) such that with a 20 MOhm bias resistor, 
the induced voltage drop is almost equal to the applied voltage.

In short, up 300V we cannot apply high enough bias to overcome the equivalent increase in leakage current.

Explanation?
This could be an effect of ions on eccentric orbits hitting something and showering the detector with electrons, leading to surface charging and the increased leakage currents.
Either we are receiving huge numbers of ions in the detector but not seeing them, or something else like above is going on.

Surface charging could/would dissipate over time and potentially bring the detectors back into a useful range. Radiation damage likely would not.
ELOG V3.1.5-fc6679b